OpenStudy (anonymous):

Simplify square root of 175x^4 (assume that all variables are non-negative real numbers)

OpenStudy (asnaseer):

\[175 = 25 * 7 = 5^2 * 7\]so:\[\sqrt{175x^4}=\sqrt{7*5^2*x^4}\]can you solve now?

OpenStudy (asnaseer):

sorry - not "solve" but "simplify"